A long-term care facility resident, Kubik, had a hip knee replacement 6 months ago and consequently developed an infection with methicillin-resistant Staphylococcus aureus and then osteomyelitis. He has a history of diabetes. 1. Kubik receives vancomycin IVPB ordered at 15 mg/kg/day. Kubik weighs 220 lb. What daily dose would the nurse administer? 2. What nursing considerations are essential when giving this medication? 3. What are the signs and symptoms and treatment the nurse would be alert for if Kubik developed red man syndrome?
Answer #1

1. ordered dos eof vancomycin=15 mg/kg/day.

Kubik weighs 220 lb=220/2.2=100 kg

Then the required dose as per the weight= 15X 100= 1500 mg will be the daily dose of the Vancomycin

2. Essential nursing considerations when giving vancomycin:

It is adrecommended to monitor the renal function
Then it is advised to avoid extravasation in the whole course of the therapy

3. The signs and symptoms about which nurse should be alert:

There are chances of "red man syndrome": sore throat, fever, reduced blood pressure and fatigue.
Answer #2

1. ordered dos eof vancomycin=15 mg/kg/day.

Kubik weighs 220 lb=220/2.2=100 kg

Then the required dose as per the weight= 15X 100= 1500 mg will be the daily dose of the Vancomycin

2. Essential nursing considerations when giving vancomycin:

It is adrecommended to monitor the renal function
Then it is advised to avoid extravasation in the whole course of the therapy

3. The signs and symptoms about which nurse should be alert:

There are chances of "red man syndrome": sore throat, fever, reduced blood pressure and fatigue.